Calculate the frequency of the two alleles showing incomplete dominance. Phenotype and population... hom fast=976 hom slow =58 heterozygous 316.
1
Expert's answer
2017-04-22T05:03:08-0400
The total number of individuals in population is 976 + 58 + 316 = 1350.
Frequency of allele for hom fast: (2*976 + 316)/2*1350 = 0.84
Frequency of allele for hom slow: 1-0.84 = 0.16